Find the coordinates of the midpoint of line HK if H (-1, 2) and K (-7, -4).

Grade:9

1 Answers

Vikas TU
3 years ago
Dear student 
H(-1,2) 
K(-7,-4)
Let P be the midpoint of HK 
So coordinate of P will be 
x Coordinate of P  = (-1-7)/2 = -4 
y Coordinate of P = (2-4)/2 = -1 
So, Coordinate of P = (-4,-1) 
Hope this will work 
Good Luck